# Enhancing Quantum Memory Lifetime with Measurement-Free Local Error Correction and Reinforcement Learning

### v1

@@ V.1 Effect of LEC circuits on error statistics
We define the average memory lifetime $T$ encoded in each code to be an average number of LEC rounds before the final recovery step calls decoding failure for the first time as discussed in Appendix B.4. Due to its relation with logical error rate, the lifetime of memory is extended more effectively by RL LECs than conventional ones. Now, based on this understanding, we motivate a parameter of the faulty LEC circuit that represents its performance in extending the lifetime of encoded memory — called effective code distance.
V.2 Idea of effective code distance
@@
First, let’s consider a case where there is no LEC, i.e. error with $p_{amb}$ introduced per qubit per round until the final recovery failure is evaluated by performing a perfect global decoding. In this case, the error rate is accumulated approximately $p_{amb}$ every round, and the average memory lifetime $T$ can be approximated as
$T\sim\frac{p_{th}}{p_{amb}},$ (8)
@@
where $p_{th}$ refers to a threshold error rate for the final decoding success evaluation step discussed in Appendix B.4. Note that $p_{th}$ depends on a type of error-correcting code: 2D toric code has $p_{th}\approx 12\%$, 2D Ising model has $p_{th}=50$%, and 4D toric code has $p_{th}\approx 4.5\%$. This behavior is because, for a large enough system size, the final recovery step can be approximated to perform successful decoding if the system’s error rate is below $p_{th}$ and failed decoding otherwise.
@@
Now, let’s consider a case with LEC: LEC rounds, which consist of error generation with rate $p_{amb}$ followed by an LEC circuit application, are repeated until the decoding failure is evaluated. First, suppose that $p_{gate}=0$. Since an LEC circuit reduces error density in the system, the LEC circuit is expected to suppress the accumulation rate of errors $p_{acc}$ in each LEC round:
$p_{acc}\propto(p_{amb})^{D_{eff}},$ (9)
@@
where the power law exponent is called effective code distance $D_{eff}$ of the LEC circuit. Combining Equation 8 and 9, we obtain that
$T\sim\frac{p_{th}}{p_{acc}}\sim(p_{amb})^{-D_{eff}}.$ (10)
@@
Thus, for the same final recovery step and $p_{amb}$ each LEC round, $T$ gets extended as $D_{eff}$ increases. Since different LEC circuits suppress error accumulation by different amounts, $D_{eff}$ is a parameter that characterizes the performance of the LEC circuit in extending the memory lifetime.
@@
In particular, $D_{eff}$ depends on the size of the uncorrectable error configuration by the given LEC circuit. Without LEC, every error in the system survives, gets passed to the next round, and contributes to the accumulation of error density. However, with LEC, primarily the “uncorrectable” error configuration by the entire LEC circuit survives and contributes to the error density accumulation. We summarize the scale of uncorrectable error by LEC (compared to logical error) in Table 2.
@@
Table 2: Number of data qubit errors to compose an effectively uncorrectable error pattern by entire LEC circuit vs. a logical error in each code. Note that $L$ denotes a system size, and $D$ denotes some constant that depends on the LEC circuit.
In Section V.1 and Appendix D, we discuss that there exist some remaining uncorrectable error patterns after applying an entire LEC circuit. We observe that in the 2D toric code, such remaining uncorrectable error patterns after applying the LEC circuit are independent of system size $L$. This is because the uncorrectable error patterns are 1D chains that are independent of $L$. However, in the 2D Ising model and 4D toric code, the uncorrectable error patterns are 2D membranes that are proportional to $L$. Thus, remaining uncorrectable error patterns after applying the LEC circuit are proportional to system size $L$ as well.
@@
Figure 9: Effect of faulty gates on effective code distance $D_{eff}$ by comparing a correction of lower-weight (thick red line) vs. higher-weight (thin red line) initial error configurations. Local decoding with the same LEC circuit over time is schematically shown for benchmark $p_{gate}=0$ and $p_{gate}\neq 0$. (a) When benchmark $p_{gate}=0$, the LEC circuit removes both of the initial error configurations at the end. (b) When benchmark $p_{gate}\neq 0$, the LEC circuit still removes an error configuration with a lower weight at the end by the local decoding process (thick orange lines with empty arrows). However, an error configuration with a higher weight “grows” over time by merging with gate errors introduced during the local decoding process (thin orange lines with filled arrows). Thus, we expect $D_{eff}$ for a given LEC circuit to decrease as benchmark $p_{gate}$ increases.
@@
Although the above discussion is for LEC circuits with perfect gates, we also define $D_{eff}$ for faulty LEC circuits as well. Note that the LEC circuit can be interpreted as a dynamical process. The initial error configuration is newly introduced to the system by ambient or gate errors. Then, the LEC circuit is applied to correct such an error pattern. When $p_{gate}=0$ as discussed above, the decoding is performed only to these initial errors as shown in Fig. 9(a). However, when $p_{gate}\neq 0$, new errors introduced after each parallel gate operation can cause the growth of initial error configuration during the decoding process as shown in Figure 9(b). Due to this effect, $D_{eff}$ gets decreased when LEC circuits consist of faulty gates. Also, these faulty gates impact the failure rate in each round as well due to this insertion of errors by gate errors during the local decoding process. Thus, we model the average memory lifetime $T$ for a given LEC circuit with faulty gates as
$T\approx C\left(\frac{1}{\alpha\times p_{amb}}\right)^{D_{eff}}.$ (11)
@@
Both $D_{eff}$ and $\alpha$ depend on the choice of error-correcting code, $p_{gate}$, and an LEC circuit. Physically, the parameter $D_{eff}$ describes how fast such errors grow over time, whereas the parameter $\alpha$ describes how many new errors (including gate errors) get introduced every LEC round. From Table 2, we expect that (1) $D_{eff}$ is independent of $L$ for 2D toric code and (2) $D_{eff}\propto L$ for 2D Ising model and 4D toric code. This intuition is confirmed in the following Section V.3. Also, we discuss the detailed intuition for $\alpha$ in Appendix F. Note that our characterization of memory lifetime by a local decoder is different from the previous literature [38], which follows the quadratic relation between lifetime and ambient error rate. This is due to the difference between gate error regimes of interest.
V.3 Fitting effective code distance from lifetime
@@
Figure 10: Fitting from log-log plot of average memory lifetime $T_{L}$ for each system size $L$ vs. ambient error rate $p_{amb}$ for fixed benchmark gate error rate $p_{gate}$ to compute effective code distance $D_{eff}$. In each code, we perform a simulation of multiple LEC cycles on 1,000 independent samples to benchmark $T_{L}$, which is the number of LEC cycles until the final perfect decoding results in a logical failure. Our computation for the error bar in average $T_{L}$ is discussed in Appendix C. We compare RL-optimized LEC circuit (blue gradation for varying $L$) vs. nearest-neighbor LEC circuit (NN LEC) in 2D toric code or Toom’s rule-based LEC circuit (Toom LEC) in 2D Ising model and in 4D toric code (grey gradation for varying $L$). RL training parameters for code and error model are as in Table 1 with $p_{gate}=1\times 10^{-3},\text{ }1\times 10^{-4},\text{ and }1\times 10^{-5}$ for 2D Ising model, 2D and 4D toric code, respectively. Note that a blue dashed line (:) corresponds to a training $p_{amb}$ for each code.
@@
Figure 11: Fitted effective code distance $D_{eff}$ for given LEC circuits vs. benchmark gate error rate $p_{gate}$ in each code. Note that each curve represents $D_{eff}$ vs. benchmark $p_{gate}$ for the same LEC circuit. Our computation for the error bar of $D_{eff}$ is discussed in Appendix C. Grey squares represent data with (a) nearest-neighbor LEC circuit (NN LEC) in 2D toric code and (b,c) Toom’s rule-based LEC circuit (Toom LEC) in 2D Ising model and 4D toric code. Blue circles represent data with RL-optimized LEC circuit trained at (a) $p_{gate}=1\times 10^{-4}$ in 2D toric code, (b) $p_{gate}=1\times 10^{-3}$ in 2D Ising model, and (c) $p_{gate}=1\times 10^{-5}$ in 4D toric code. Orange diamonds represent data with RL-optimized LEC circuit trained at (a) $p_{gate}=2\times 10^{-4}$ in 2D toric code, (b) $p_{gate}=5\times 10^{-3}$ in 2D Ising model, and (c) $p_{gate}=2\times 10^{-5}$ in 4D toric code.
@@
In Equation 11, the parameter $C$, $\alpha$, and $D_{eff}$ can be determined by fitting the data between average memory lifetime $T_{L}$ vs. ambient error rate $p_{amb}$ as shown in Fig. 10. To achieve this fitting, we benchmark the average memory lifetime $T_{L}$ in each error-correcting code with varying input parameters: LEC circuit, $L$, $p_{amb}$, and $p_{gate}$. In the 2D toric code, we compare the NN LEC and RL LEC trained with different $p_{gate}$. In the 2D Ising model and 4D toric code, we compare the Toom LEC and RL LEC trained with different $p_{gate}$. $T_{L}$ for each of these LEC circuits is benchmarked with $L\in\{8,12,16\}$ for 2D toric code and 2D Ising model and with $L\in\{4,6\}$ for 4D toric code.
Based on the data obtained from these benchmark simulations, we perform a fitting to find $D_{eff}$ in each error-correcting code. Note that the fitting lines in Fig. 10 are obtained from the following fitting models.
@@
In 2D toric code, $T$ is approximately independent of $L$ as confirmed in Fig. 10 — with a caveat as discussed in Appendix E. Thus, we fit the following model between $T$ obtained by simulation and $p_{amb}$:
$\log(T)=-{\color[rgb]{1,0,0}D_{eff}}\times\log(p_{amb})+{\color[rgb]{1,0,0}k},$ (12)
where $k$ is some constant that corresponds to $\log(C)-D_{eff}\log(\alpha)$ from parameters in Equation 11.
@@
In 2D Ising model and 4D toric code, $D_{eff}$ is expected to be proportional to $L$. Thus, we fit the following model between $T_{L}$ (average memory lifetime obtained by simulating a code with system size $L$) and $p_{amb}$:
$\log(T_{L})=-{\color[rgb]{1,0,0}\frac{D_{eff}}{L}}\times L\times(\log(p_{amb})+{\color[rgb]{1,0,0}k_{1}})+{\color[rgb]{1,0,0}k_{2}},$ (13)
where $k_{1}$ and $k_{2}$ correspond to $\log(\alpha)$ and $\log(C)$, respectively, from parameters in Equation 11.
@@
With a fixed LEC circuit, $L$, and benchmark $p_{gate}$, let each dataset be the pairs of benchmark $p_{amb}$ and average lifetime. Then, to perform a global fitting based on Equation 12 and 13, we use the symfit python package [90]. Parameters $(D_{eff},k)$ for the 2D toric code and $(D_{eff},k_{1},k_{2})$ for the 2D Ising model and 4D toric code were shared by datasets varying $L$ with the same LEC circuit and benchmark $p_{gate}$. Although we are mainly interested in $D_{eff}$ in each code, we also discuss $k_{1}$ fitting result based on our intuition of $\alpha$ in Appendix F.
@@
After the fitting, we plot $D_{eff}$ of each LEC circuit vs. benchmark $p_{gate}$ as shown in Fig. 11. Let $p_{gate,c}$ for RL LEC be a critical gate error rate such that RL LEC has higher $D_{eff}$ compared to NN LEC (if 2D toric code) or Toom LEC (if 2D Ising model or 4D toric code). In every error-correcting code, we observe that the RL LEC trained with lower $p_{gate}$ shows higher fitted $D_{eff}$ in small $p_{gate}$ regime but smaller $p_{c}$ compared to the LEC circuit trained with higher $p_{gate}$. For the LEC circuits that are trained with small $p_{gate}$ as shown in Fig. 11, we additionally find that $p_{gate,c}$ is larger than the training $p_{gate}$. Note that this property does not apply to LEC circuits trained with large $p_{gate}$ as discussed in Appendix G. Also, in the 2D Ising model, the LEC circuit only with Toom actions shows fitted $D_{eff}/L\approx 1.0$. This is because the minimum uncorrectable error configuration by the Toom LEC circuit consists of $L$ number of data qubit errors as shown in Fig. 4(b). Note that this observation and its explanation are consistent with the intuition in Section V.2, especially Table 2.
@@ VI Application of LEC Scheme
In this section, we first discuss the experimental implementation of the LEC scheme in Section VI.1. Then, we provide two specific examples of how this optimized LEC scheme could be used in various quantum information processing tasks. Section VI.2 explores how RL-optimized LEC circuits can be used to reduce the number of mid-circuit readouts required to preserve a quantum memory in 2D toric code. On the other hand, Section VI.3 discusses the dissipative preparation of topological phases with a given Hamiltonian.
VI.1 Experimental implementation of LEC
VI.2 Reducing rate of mid-circuit readouts
Figure 12: Original vs. Hybrid decoding protocol to preserve 2D toric code memory. (a) At each unit time, new errors with error rate $p_{unit}=0.1\%$ are introduced to data qubits. Throughout total physical time $T$, global decoding with faulty readouts and 3D minimum-weight perfect matching (MWPM) algorithm (blue, solid lines) is performed regularly. Original and hybrid protocol performs it $N_{MWPM}^{(Original)}$ and $N_{MWPM}^{(Hybrid)}$ times, respectively. However, hybrid decoding protocol applies LEC circuit (orange, dotted lines) $N_{LEC}$ times between global decoding applications. After $T$ passes, we perform global decoding with perfect readouts to evaluate the final logical error rate $P_{L}$. (b) Method of finding minimum $N_{MWPM}^{(Original)}$ and $N_{MWPM}^{(Hybrid)}$ to achieve $P_{L}<15$% for fixed $T=2000$ with 5,000 samples. Our computation for the error bar of $P_{L}$ is discussed in Appendix C. (c) Minimum $N_{MWPM}^{(Original)}$ and $N_{MWPM}^{(Hybrid)}$ vs. $p_{unit}\times T$ for given multi-qubit gate fidelities and $P_{L}<15$%.
In Section V, our results show that LEC circuits can be used to remove errors from encoded qubits and to extend the memory lifetime. Then, it is reasonable to ask whether such LEC circuits can be beneficial to extend memory lifetime in practice. We focus on a 2D toric code to investigate the practical application.

### v2

@@ V.1 Idea of effective code distance
We define the average memory lifetime $T$ encoded in each code to be an average number of LEC rounds before the final recovery step calls decoding failure for the first time. We now introduce a parameter that represents the faulty LEC circuit’s performance in extending the memory lifetime — called effective code distance.
@@
First, let’s consider a case without LEC, i.e. error with $p_{\rm{amb}}$ introduced per qubit per round until the final recovery failure after multiple rounds is evaluated by performing a perfect global decoding. In this case, the error rate is accumulated approximately $p_{\rm{amb}}$ every round, and the average memory lifetime $T$ can be approximated as
$T\sim\frac{p_{\rm{th}}}{p_{\rm{amb}}},$ (2)
@@
where $p_{\rm{th}}$ refers to a threshold error rate for the final decoding success evaluation step. Note that $p_{\rm{th}}$ depends on a type of code: 2D toric code has $p_{\rm{th}}\approx 12\%$, 2D Ising model has $p_{\rm{th}}=50$%, and 4D toric code has $p_{\rm{th}}\approx 4.5\%$. This behavior is because, for a large enough system size, the final recovery step can be approximated to perform successful decoding if the system’s error rate is below $p_{\rm{th}}$ and failed decoding otherwise.
@@
Now, let’s consider a case with LEC: we repeat error generation with rate $p_{\rm{amb}}$ followed by an LEC circuit application until the decoding failure is evaluated. First, suppose that gates are perfect, i.e., $p_{\rm{gate}}=0$. Since an LEC circuit reduces error density in the system, the LEC circuit is expected to suppress the accumulated rate of errors $p_{\rm{acc}}$ in each LEC round:
$p_{\rm{acc}}\propto(p_{\rm{amb}})^{D_{\rm{eff}}},$ (3)
@@
where the power law exponent is called effective code distance $D_{\rm{eff}}$ of the LEC circuit. Combining Equation 2 and 3, we obtain that
$T\sim\frac{p_{\rm{th}}}{p_{\rm{acc}}}\sim(p_{\rm{amb}})^{-D_{\rm{eff}}}.$ (4)
@@
Thus, for the same final recovery step and $p_{\rm{amb}}$ each LEC round, $T$ gets extended as $D_{\rm{eff}}$ increases. Since different LEC circuits suppress error accumulation by different amounts, $D_{\rm{eff}}$ characterizes the performance of the LEC circuit in extending the memory lifetime.
@@
In particular, $D_{\rm{eff}}$ depends on the size of the uncorrectable error configuration by the given LEC circuit. Without LEC, every error in the system survives, gets passed to the next round, and contributes to the accumulation of error density. However, with LEC, primarily the “uncorrectable” error configuration by the entire LEC circuit survives and contributes to the error density accumulation. We summarize the scale of uncorrectable error by LEC compared to logical error in Table 1.
@@
Table 1: Number of data qubit errors to compose an effectively uncorrectable error pattern by entire LEC circuit vs. a logical error in each code. $L$ denotes a system size, and $D$ denotes some constant that depends on the LEC circuit.
@@
Figure 8: Effect of faulty gates on effective code distance by comparing a correction of lower-weight (thick red line) vs. higher-weight (thin red line) initial error configurations. Local decoding with the same LEC circuit over time is schematically shown for perfect vs. faulty gates. (a) When gates are perfect, the LEC circuit removes both of the initial error configurations at the end. (b) When gates are faulty, the LEC circuit still removes an error configuration with a lower weight at the end by the local decoding process (thick orange lines with empty arrows). However, an error configuration with a higher weight “grows” over time by merging with gate errors introduced during the local decoding process (thin orange lines with filled arrows). Thus, we expect the effective code distance for a given LEC circuit to decrease as the circuit’s gate error rate increases.
@@
Although the above discussion is for LEC circuits with perfect gates, we also define $D_{\rm{eff}}$ for faulty LEC circuits as well (i.e., $p_{\rm{gate}}\neq 0$). Note that the LEC circuit can be interpreted as a dynamical process. The initial error configuration is newly introduced to the system by ambient or gate errors. Then, the LEC circuit is applied to correct such an error pattern. When gates are perfect, the decoding is performed only to these initial errors as shown in Fig. 8(a). However, when gates are faulty, new errors introduced after each parallel gate operation can cause the growth of initial error configuration during the decoding process as shown in Figure 8(b). Due to this effect, $D_{\rm{eff}}$ gets decreased when LEC circuits consist of faulty gates. Also, these faulty gates impact the failure rate in each round as well due to this insertion of errors by gate errors during the local decoding process. Thus, we model the average memory lifetime $T$ for a given LEC circuit with faulty gates as
$T\approx C\left(\frac{1}{\alpha\times p_{\rm{amb}}}\right)^{D_{\rm{eff}}}.$ (5)
@@
Both $D_{\rm{eff}}$ and $\alpha$ depend on the choice of error-correcting code, $p_{\rm{gate}}$, and an LEC circuit. Physically, the parameter $D_{\rm{eff}}$ describes how fast such errors grow over time, whereas the parameter $\alpha$ describes how many new errors (including gate errors) get introduced every LEC round. From Table 1, we expect that (1) $D_{\rm{eff}}$ is independent of $L$ for 2D toric code and (2) $D_{\rm{eff}}\propto L$ for 2D Ising model and 4D toric code. Our characterization of memory lifetime by a local decoder is different from the previous literature [38], which follows the quadratic relation between lifetime and ambient error rate. This is due to the difference between gate error regimes of interest.
V.2 Fitting effective code distance
@@
Figure 9: Fitting from log-log plot of average memory lifetime $T_{L}$ for each system size $L$ vs. ambient error rate $p_{\rm{amb}}$ for fixed benchmark gate error rate $p_{\rm{gate}}$ to compute effective code distance. In each code, we perform a simulation of multiple LEC cycles on 1000 independent samples to benchmark $T_{L}$, which is the number of LEC cycles until the final perfect decoding results in a logical failure (see Appendix D for the error bar). We compare RL-optimized LEC circuit (blue gradation for varying $L$) vs. nearest-neighbor LEC circuit (NN LEC) in 2D toric code or Toom’s rule-based LEC circuit (Toom LEC) in 2D Ising model and in 4D toric code (grey gradation for varying $L$). We train the RL model with $p_{\rm{gate}}=1\times 10^{-3},\text{ }1\times 10^{-4},\text{ and }1\times 10^{-5}$ for 2D Ising model, 2D and 4D toric code, respectively. Note that a blue dashed line (:) corresponds to training $p_{\rm{amb}}$ for each code.
@@
Figure 10: Fitted effective code distance $D_{\rm{eff}}$ for given LEC circuits vs. benchmark gate error rate $p_{\rm{gate}}$ in each code (see Appendix D for the error bar). Grey squares represent data with (a) nearest-neighbor LEC circuit (NN LEC) in 2D toric code and (b,c) Toom’s rule-based LEC circuit (Toom LEC) in 2D Ising model and 4D toric code. Blue circles (or orange diamond) represent data with RL-optimized LEC circuit trained at $p_{\rm{gate}}$ with value (a) $1\times 10^{-4}$ (or $2\times 10^{-4}$) in 2D toric code, (b) $1\times 10^{-3}$ (or $5\times 10^{-3}$) in 2D Ising model, and (c) $1\times 10^{-5}$ (or $2\times 10^{-5}$) in 4D toric code.
@@
In Equation 5, the parameter $C$, $\alpha$, and $D_{\rm{eff}}$ can be determined by fitting the data between average memory lifetime $T_{L}$ vs. ambient error rate $p_{\rm{amb}}$ as shown in Fig. 9. To achieve this fitting, we benchmark the average memory lifetime $T_{L}$ in each error-correcting code with varying input parameters. Then, we compare the RL LEC trained with different $p_{\rm{gate}}$ and conventional LEC.
@@
In 2D toric code, $T$ is approximately independent of $L$ (see Appendix E.2 for a caveat). Thus, from Equation 5, we fit the following model between $T$ and $p_{\rm{amb}}$:
$\log_{10}(T)=-{\color[rgb]{1,0,0}D_{\rm{eff}}}\times\log_{10}(p_{\rm{amb}})+{\color[rgb]{1,0,0}k},$ (6)
@@
where $k=\log_{10}C-D_{\rm{eff}}\times\log_{10}\alpha$. In 2D Ising model and 4D toric code, $D_{\rm{eff}}$ is expected to be proportional to $L$. Thus, from Equation 5, we fit the following model between $T_{L}$ (average memory lifetime obtained by simulating a code with system size $L$) and $p_{\rm{amb}}$:
$\log_{10}(T_{L})=-{\color[rgb]{1,0,0}\frac{D_{\rm{eff}}}{L}}\times L\times(\log_{10}(p_{\rm{amb}})+{\color[rgb]{1,0,0}k_{1}})+{\color[rgb]{1,0,0}k_{2}},$ (7)
where $k_{1}$ and $k_{2}$ are $\log_{10}\alpha$ and $\log_{10}C$, respectively.
@@
With a fixed LEC circuit, $L$, and benchmark $p_{\rm{gate}}$, let each dataset be the pairs of benchmark $p_{\rm{amb}}$ and average lifetime. Then, to perform a global fitting based on Equation 6 and 7, we use the symfit python package [96]. Parameters $(D_{\rm{eff}},k)$ for the 2D toric code and $(D_{\rm{eff}},k_{1},k_{2})$ for the 2D Ising model and 4D toric code were shared by datasets varying $L$ with the same LEC circuit and benchmark $p_{\rm{gate}}$.
@@
After the fitting, we plot $D_{\rm{eff}}$ of each LEC circuit vs. benchmark $p_{\rm{gate}}$ as shown in Fig. 10 (see Appendix E.3 for $k_{1}$ fitting). Let $p_{\rm{gate,c}}$ for RL LEC be a critical gate error rate such that RL LEC has higher $D_{\rm{eff}}$ compared to NN LEC (if 2D toric code) or Toom LEC (if 2D Ising model or 4D toric code). In every error-correcting code, we observe that the RL LEC trained with lower $p_{\rm{gate}}$ shows higher fitted $D_{\rm{eff}}$ in small $p_{\rm{gate}}$ regime but smaller $p_{\rm{c}}$ compared to the LEC circuit trained with higher $p_{\rm{gate}}$. For the LEC circuits that are trained with small $p_{\rm{gate}}$ as shown in Fig. 10, we additionally find that $p_{\rm{gate,c}}$ is larger than the training $p_{\rm{gate}}$. Note that this property does not apply to LEC circuits trained with large $p_{\rm{gate}}$ (see Appendix E.4). Also, in the 2D Ising model, the LEC circuit only with Toom actions shows fitted $D_{\rm{eff}}/L\approx 1.0$. This is because the minimum uncorrectable error configuration by the Toom LEC circuit consists of $L$ number of data qubit errors as shown in Fig. 4(b). Note that this observation and its explanation are consistent with the intuition from Table 1.
@@ VI Application of LEC Scheme
In this section, we provide two specific examples of how this optimized LEC scheme could be used in various quantum information processing tasks. Section VI.1 explores how RL-optimized LEC circuits can be used to reduce the number of mid-circuit readouts required to preserve a quantum memory in 2D toric code. On the other hand, Section VI.2 discusses the dissipative preparation of topological phases with a given Hamiltonian.
VI.1 Reducing rate of mid-circuit readouts